A charge of +6.0µC experiences a force of 2.0 mN in the +x-direction at a certain point in space. (a) What was the electric field there before the charge was placed there? (b) Describe the force a - 2.0µC charge would experience if it were used in place of the +6.0µC charge.


1
Expert's answer
2022-03-25T15:02:38-0400

(a) The electric field is directed toward positive x-axis. The magnitude of field

"E=\frac{F}{q}=\frac{2.0*10^{-3}}{6.0*10^{-6}}=333\:\rm N/C"

(b) The force on the negative charge is directed opposite to the field. The magnitude of force

"F=qE\\
=2.0*10^{-6}*333=6.7*10^{-4}\;\rm N"